OpenStudy (anonymous):

Solve the exponential inequality. 3^3+x > 729......... for your information 3+x is an exponent

OpenStudy (amistre64):

since the exponent is a grouping, in the future you may just want to use ( ) to wrap it up in 3^(3+x) > 729

OpenStudy (anonymous):

OK

OpenStudy (amistre64):

exponents have the property that: exp(a+b) = exp(a) * exp(b)

OpenStudy (amistre64):

3^(3+x) > 729 3^(3) * 3^(x) > 729 3^(x) > 729/27 is a good start

OpenStudy (anonymous):

then you'd have 3^(x)>27

OpenStudy (amistre64):

correct, which should be small enough to do without a calculator now

OpenStudy (anonymous):

so would x equal 9

OpenStudy (amistre64):

3^1 = 3 3^2 = 9 3^3 = 27 3^x > 27 3^x > 3^3 x > 3

OpenStudy (anonymous):

so the reason we got 3 is because?????

OpenStudy (amistre64):

because 27 is equal to 3^3 3^x is greater than 3^3 we are comparing the x values ... if x=3, then 3^3 = 3^3 ... we want it to be greater than so x has to be greater than 3

OpenStudy (amistre64):

we could log it out if you wanted a more mathical approach 3^x > 3^3 log(3^x) > log(3^3) x log(3) > 3 log(3) x > 3

OpenStudy (anonymous):

THANK YOU i have another question that im half way done with idk if you want to help though

OpenStudy (amistre64):

if its small enough sure

OpenStudy (anonymous):

ok 5^(2x+2) < 625 and i already have 5^(x+2)<25 whats the next step

OpenStudy (amistre64):

split the exponent and divide out the constant

OpenStudy (amistre64):

only your setup is flawed

OpenStudy (amistre64):

5^(2x+2) < 625 5^[2(x+1)] < 625 25^(x+1) < 625

OpenStudy (amistre64):

25^(x+1) < 625 25^(x)+ * 25^(1) < 625 25^(x) < 625/25

OpenStudy (amistre64):

second line has a spurious little + hanging in there ... just ignore it lol

OpenStudy (amistre64):

25^(x+1) < 625 25^(x)* 25^(1) < 625 ... thats better 25^(x) < 625/25

OpenStudy (anonymous):

lol i got 25^x < 25 the answer would be x<1

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

yea buddy lol thanx my friend that is all

OpenStudy (amistre64):

youre welcome ... good luck :)